Minutes of the Trigger Operation Meeting (19th Dec, 2006)
Attendees: Debbie Bard, Al Eisner, Selina Li, Su Dong.
- Spare BLTs had their old PROMs replaced and were tested to be working at the teststand. Board #3505 is down in IR2 spares cabinet and board #2210 is in the lab. #2210 had LEDs that light up backwards, which is odd.
- Cosmics run 69139 looks fine from fastmon plots, but L1TNtuple isn't working at the moment (probably because of computing outages) so can't check it all yet. The IFT timing looks alright, but can't tell without looking at L1TNtuple data! Should check with IFR people if the switch was flipped to use the correct timing.
- LiveFastMon plots have old labels! Should be updated.
- Selina is kindly taking the pager while Debbie is away over christmas.
- Mark has spoken to Dan about the EMT MIP plots and efficiencies. There was a question about why there's entries at low energy (~0.02GeV), but that's to be expected with random interactions.
- The sliceing and presenting of Bhabha plots should be kept in as a part of post-run processing. The changes are known and should be done sometime soon.
- Question of possibly increasing the number of background triggers, so there's more in the simualtion. This is more of a question for computing people as to whether there's space to store this.
This page is created by Debbie Bard (firstname.lastname@example.org); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.